# Galactic echoes

### v1

@@ Abstract
Galaxy: kinematics and dynamics – methods: analytical – plasma physics
@@ 1 Introduction
The Gaia ‘Snail’ (Antoja et al., 2018) is a one-armed spiral structure in the density of stars in the Solar neighborhood when viewed in vertical phase-space $(z,v_{z})$. Here, $z$ is the vertical position of a star normal to the galactic mid-plane and $v_{z}$ is its vertical velocity. The existence of the Snail demonstrates that the Milky Way is not in phase-mixed equilibrium. In fact, Snail-like structures manifest not only in stellar density, but also in stellar labels like age and metallicity (Frankel et al., 2024). Snails almost certainly result from the incomplete phase-mixing of the stars’ distribution function (DF) following dynamical perturbations of some kind, but what precisely were/are these dynamical perturbations remains an unanswered question (see the review by Hunt & Vasiliev 2025 and references therein).
@@ 2.2.1 General solution
×$\times\mathrm{i}\sum_{n^{\prime}}\bigg{[}n^{\prime}\frac{\partial\hat{\Phi}_{1,n-n^{\prime}}(J,t^{\prime})}{\partial J}-(n-n^{\prime})\hat{\Phi}_{1,n-n^{\prime}}(J,t^{\prime})\frac{\partial}{\partial J}\bigg{]}$
×$\times\int_{0}^{t^{\prime}}\mathrm{d}t^{\prime\prime}\,\mathrm{e}^{-\mathrm{i}n^{\prime}\Omega(t^{\prime}-t^{\prime\prime})}\mathrm{i}n^{\prime}\frac{{\partial}f_{0}}{{\partial}J}\hat{\Phi}_{1,n^{\prime}}(J,t^{\prime\prime}).$ (14)
@@
×$\times\bigg{[}n^{\prime}{\partial}_{J}\hat{\Phi}_{1,n-n^{\prime}}(J,t^{\prime})-(n-n^{\prime})\hat{\Phi}_{1,n-n^{\prime}}(J,t^{\prime})D_{J}(t^{\prime})\bigg{]}$
×$\times\frac{{\partial}f_{0}}{{\partial}J}\hat{\Phi}_{1,n^{\prime}}(J,t^{\prime\prime})\,\mathrm{e}^{-\mathrm{i}\Omega[nt-(n-n^{\prime})t^{\prime}-n^{\prime}t^{\prime\prime}]},$ (16)
@@ 2.2.2 Two impulsive kicks
×$\times\bigg{\{}\bigg{[}n^{\prime}{\partial}_{J}\psi_{1}-(n-n^{\prime})\psi_{1}D_{J}(t^{\prime})\bigg{]}\delta_{n-n^{\prime}}^{s_{a}n_{1}}\delta(t^{\prime}-t_{1})+(1\leftrightarrow 2)\bigg{\}}$
×$\times\frac{{\partial}f_{0}}{{\partial}J}\bigg{[}\psi_{1}\delta_{n^{\prime}}^{s_{b}n_{1}}\delta(t^{\prime\prime}-t_{1})+(1\leftrightarrow 2)\bigg{]}\,\mathrm{e}^{-\mathrm{i}\Omega[nt-(n-n^{\prime})t^{\prime}-n^{\prime}t^{\prime\prime}]}.$
@@
×$\times\bigg{\{}n_{1}\psi_{1}{\partial}_{J}\psi_{2}\frac{{\partial}f_{0}}{{\partial}J}+n_{2}\psi_{1}\psi_{2}\frac{{\partial}^{2}f_{0}}{{\partial}J^{2}}+n_{2}\psi_{2}{\partial}_{J}\psi_{1}\frac{{\partial}f_{0}}{{\partial}J}$
@@
×$\times\bigg{[}\big{(}n_{1}\psi_{1}{\partial}_{J}\psi_{2}+n_{2}\psi_{2}{\partial}_{J}\psi_{1}\big{)}\frac{{\partial}f_{0}}{{\partial}J}+n_{2}\psi_{1}\psi_{2}\frac{{\partial}^{2}f_{0}}{{\partial}J^{2}}$
@@ 2.3 Limitations and extensions
Third, we have ignored the self-gravity of the perturbed stellar distribution. As shown by Darling & Widrow (2019) and Widrow (2023), self-gravity may strengthen the amplitude of the Snail-like response to an impulsive perturbation, and extend the lifetime of a Snail well beyond that expected due to phase mixing alone. It is unclear what impact self-gravity would have on galactic echoes. On the one hand, we expect that if self-gravity leads to stronger first-order responses, it should lead to stronger second- and higher-order effects like echoes too. On the other hand, the $t_{2}-t_{1}$ dependence of the echo term arises precisely because nonlinear couplings feed off phase-space gradients, and under normal phase mixing these get sharper as $t_{2}-t_{1}$ increases. Since self-gravity will reduce the rate of phase mixing, it may make echoes weaker, not stronger. In a homogeneous plasma, one can perform a self-consistent calculation (Gould et al., 1967; O’Neil & Gould, 1968), which predicts that collective effects amplify the echo while slightly modifying its growth and decay rates. However, to perform such a self-consistent calculation here would incur a significant mathematical penalty: not only would we need to Laplace transform from the time-domain into the complex frequency domain (Gould et al., 1967), but also, since we are dealing here with an inhomogeneous system, we would need to project our fluctuations onto a biorthogonal potential-density basis (e.g., Kalnajs, 1976; Hamilton & Fouvry, 2024). Alternatively, we could forego both of these transforms and just solve the linearized equations numerically (Widrow, 2023). Since our aim here is to demonstrate the existence of an echo phenomenon rather than perform detailed modeling of the Solar neighborhood, we will not investigate the role of self-gravity further in this paper.
@@ 3.1 Model
We subject our disc to two impulsive perturbations caused by passing satellites or dark-matter subhalos. According to the linear response analysis by Banik et al. (2022, 2023), such encounters can excite both dipole ($n=1$) and quadrupole ($n=2$) modes in the vertical phase-space distribution of stars. The relative amplitudes of these modes depend critically on the parameters of the disc-satellite encounter, in particular, the encounter speed, $v_{\rm p}$, and the angle at which the satellite crosses the disc, $\phi_{\rm p}$, where $\phi_{\rm p}=0$ corresponds to a perpendicular passage through the disc.
To illustrate how two perturbations with different wave numbers can generate an echo, we adopt a simple scenario in which the first satellite imparts a uniform kick $\Delta v_{1}(z)=\Delta V_{1}$ to stars at $t=t_{1}$, exciting primarily a dipole-like ($n_{1}=1$) perturbation. As a rough estimate, a satellite of mass $M_{\rm p}=4\times 10^{9}{\rm M}_{\odot}$, velocity $v_{\rm p}=300\,{\rm kpc}\,{\rm Gyr}^{-1}$, and impact parameter $b=10\,{\rm kpc}$ would impart a kick of order $\Delta V_{1}\sim{\rm G}M_{\rm p}/bv_{\rm p}\sim 6\,{\rm kpc}\,{\rm Gyr}^{-1}$ (see Binney & Schönrich 2018 for a more elaborate modelling). Later, at $t=t_{2}$, we assume that the disc encounters another satellite/subhalo which crosses the vicinity of the Solar neighborhood at a grazing angle, thereby inducing a quadrupole-like ($n_{2}=2$) perturbation. We determine the $z$ dependence of the second kick by assuming that the satellite is a Plummer sphere, crossing the disc at $\phi_{\rm p}=\pi/2$. The net kick imparted to stars in the impulsive limit is then (see Appendix A)
@@
and $a$ is the core radius. In our fiducial setup, we adopt $t_{1}=0.2\,{\rm Gyr}$, $t_{2}=0.8\,{\rm Gyr}$, $\Delta V_{1}=6\,{\rm kpc}\,{\rm Gyr}^{-1}$, $\Delta V_{2}=6\,{\rm kpc}\,{\rm Gyr}^{-1}$ and $a=0.3\,{\rm kpc}$. The corresponding mass of the second satellite/subhalo is $M_{\rm p}\sim 1.2\times 10^{8}{\rm M}_{\odot}$, although this estimate can be much larger if the satellite passes at a non-zero distance from the Solar neighborhood.
@@ 3.2 Emergence of echo in the disc’s vertical phase space
The first uniform kick (blue) shifts the entire distribution towards higher velocity, resulting in an overdensity at $\theta=0$ and underdensity at $\theta=\pm\pi$. This dipole ($n_{1}=1$) perturbation subsequently shears into a one-armed phase spiral in the $(z,v)$ plane. In the $(\theta,\Omega)$ plane, this spiral appears as an inclined, rectilinear stripe. The second antisymmetric kick (green) introduces an quadrupole ($n_{2}=2$) perturbation, which evolves into a two-armed phase spiral. In addition to generating its own response, the second kick also perturbs the pre-existing one-armed phase spiral. This second-order perturbation results in a new anti-wound spiral pattern, which appears as a linear pattern with negative inclination in the $(\theta,\Omega)$ plane. As time progresses, this pattern unwinds (phase ‘unmixes’), and at the predicted echo time (red), its phase fully aligns, giving rise to an echo. The echo subsequently phase mixes and forms a new one-armed phase spiral pattern.
The phase alignment at $t_{\rm echo}$ implies that the echo temporarily causes a fluctuation in macroscopic (integrated) quantities. To illustrate this, we plot in the top panel of Fig. 2 the time evolution of the mean height of the disc, $\langle z\rangle=\int{\rm d}z{\rm d}vf(z,v,t)z$. Following the first kick, the mean height undergoes a rapid oscillation, whose amplitude gradually decays due to phase mixing. The second kick induces a density perturbation symmetric about the disc mid-plane and therefore leaves the mean height unaffected. Later, as the echo unwinds and rewinds, the disc’s mean height begins to oscillate again spontaneously, despite the lack of any external forcing. This demonstrates how the memory of earlier perturbations can be preserved in phase space and reappear later in macroscopic space. The remaining panels of Fig. 2 show a similar (though less dramatic) echo effect in the measurement of the disc’s ‘thickness’ $[\langle z^{2}\rangle-\langle z\rangle^{2}]^{1/2}$ and in $\langle z^{2}\rangle$ alone. The effect is less pronounced because the echo has most of its power at $n=\pm 1$, which does not contribute to even moments of $z$.
Figure 2: Time evolution of the mean height of the disc $\langle z\rangle$, its thickness $[\langle z^{2}\rangle-\langle z\rangle^{2}]^{1/2}$, and the rms vertical coordinate $\langle z^{2}\rangle$, all in the absence of diffusion. The mean height undergoes damped oscillation after the first uniform kick at $t_{1}$ due to phase mixing. The second antisymmetric kick at $t_{2}$ does not directly affect the mean height, although it distorts the pre-existing one-armed phase spiral, generating a renewed oscillation (an echo) near $t_{\rm echo}$. The echo signals in the other panels are weaker (note the shrunken vertical axis).
@@ 4.1 Implications for the Gaia Snail
[…]
Suppose, quite reasonably, that the first kick was provided by the Sagittarius dwarf galaxy, which is expected to primarily excite an $n_{1}=1$ mode (e.g., Banik et al., 2023). Fig. 5 then implies that, for the echo to account for the Gaia Snail, the second satellite/subhalo must have passed near the Solar neighborhood at a glancing angle $\sim 0.4\,{\rm Gyr}$ after the passage of Sagittarius and delivered an antisymmetric kick of strength $\Delta V_{2}\gtrsim 6\,{\rm kpc}\,{\rm Gyr}^{-1}$, which requires a mass of $M_{\rm p}\gtrsim 10^{8}{\rm M}_{\odot}$3.1). However, there are no known luminous satellites of such mass near the Galactic disc apart from Sagittarius (see Banik et al. 2022 and references therein). It is possible that the second kick was caused by an unseen dark subhalo, although such a massive subhalo would likely carry a detectable number of baryons. Moreover, several $N$-body simulations have shown that it is difficult to reproduce the Snail even by Sagittarius alone unless its final mass is much higher than is observed (e.g., Bland-Hawthorn & Tepper-García, 2021; Bennett et al., 2022; Asano et al., 2025). Since the (second-order) echo is inherently weaker than the (first-order) response probed in those studies, we expect that echoes would face even greater difficulty in reproducing the observed amplitude.
A more realistic scenario would be that weak echoes from perturbations by Sagittarius and other satellites/subhalos are continually overlapping with the direct first-order responses, forming phase spirals with complicated variation in amplitude and winding time across both the vertical and horizontal phase space. The residuals between the data and models based on a single encounter indeed show coherent structures, indicating the presence of multiple overlapping phase-spirals (Frankel et al., 2023). Disentangling the precise origin(s) of the Snails remains a challenging task and lies far beyond the scope of this paper. On this front, various stellar labels like age and chemical abundances, which also exhibit Snail-like patterns, may offer valuable clues, as different labels have different gradients in vertical action and angular momentum, and can thus help constrain the nature of the perturbations (Frankel et al., 2024).

### v2

@@ Abstract
Galaxy: kinematics and dynamics – Galaxy: evolution – methods: analytical
@@ 1 Introduction
The Gaia ‘Snail’ (Antoja et al., 2018) is a one-armed spiral structure in the density of stars in the Solar neighborhood when viewed in vertical phase-space $(z,v_{z})$. Here, $z$ is the vertical position of a star normal to the galactic mid-plane and $v_{z}$ is its vertical velocity. The existence of the Snail demonstrates that the Milky Way is not in phase-mixed equilibrium. In fact, Snail-like structures manifest not only in stellar density, but also in stellar labels like age and metallicity (Frankel et al., 2025). Snails almost certainly result from the incomplete phase-mixing of the stars’ distribution function (DF) following dynamical perturbations of some kind, but what precisely were/are these dynamical perturbations remains an unanswered question (see the review by Hunt & Vasiliev 2025 and references therein).
@@ 2.2.1 General solution
$\times\mathrm{i}\sum_{n^{\prime}}\bigg[n^{\prime}\frac{\partial\hat{\Phi}_{1,n-n^{\prime}}(J,t^{\prime})}{\partial J}-(n-n^{\prime})\hat{\Phi}_{1,n-n^{\prime}}(J,t^{\prime})\frac{\partial}{\partial J}\bigg]$
$\times\int_{0}^{t^{\prime}}\mathrm{d}t^{\prime\prime}\,\mathrm{e}^{-\mathrm{i}n^{\prime}\Omega(t^{\prime}-t^{\prime\prime})}\mathrm{i}n^{\prime}\frac{{\partial}f_{0}}{{\partial}J}\hat{\Phi}_{1,n^{\prime}}(J,t^{\prime\prime}).$ (14)
@@
$\times\bigg[n^{\prime}{\partial}_{J}\hat{\Phi}_{1,n-n^{\prime}}(J,t^{\prime})-(n-n^{\prime})\hat{\Phi}_{1,n-n^{\prime}}(J,t^{\prime})D_{J}(t^{\prime})\bigg]$
$\times\frac{{\partial}f_{0}}{{\partial}J}\hat{\Phi}_{1,n^{\prime}}(J,t^{\prime\prime})\,\mathrm{e}^{-\mathrm{i}\Omega[nt-(n-n^{\prime})t^{\prime}-n^{\prime}t^{\prime\prime}]},$ (16)
@@ 2.2.2 Two impulsive kicks
$\times\bigg\{\bigg[n^{\prime}{\partial}_{J}\psi_{1}-(n-n^{\prime})\psi_{1}D_{J}(t^{\prime})\bigg]\delta_{n-n^{\prime}}^{s_{a}n_{1}}\delta(t^{\prime}-t_{1})+(1\leftrightarrow 2)\bigg\}$
$\times\frac{{\partial}f_{0}}{{\partial}J}\bigg[\psi_{1}\delta_{n^{\prime}}^{s_{b}n_{1}}\delta(t^{\prime\prime}-t_{1})+(1\leftrightarrow 2)\bigg]\,\mathrm{e}^{-\mathrm{i}\Omega[nt-(n-n^{\prime})t^{\prime}-n^{\prime}t^{\prime\prime}]}.$
@@
$\times\bigg\{n_{1}\psi_{1}{\partial}_{J}\psi_{2}\frac{{\partial}f_{0}}{{\partial}J}+n_{2}\psi_{1}\psi_{2}\frac{{\partial}^{2}f_{0}}{{\partial}J^{2}}+n_{2}\psi_{2}{\partial}_{J}\psi_{1}\frac{{\partial}f_{0}}{{\partial}J}$
@@
$\times\bigg[\big(n_{1}\psi_{1}{\partial}_{J}\psi_{2}+n_{2}\psi_{2}{\partial}_{J}\psi_{1}\big)\frac{{\partial}f_{0}}{{\partial}J}+n_{2}\psi_{1}\psi_{2}\frac{{\partial}^{2}f_{0}}{{\partial}J^{2}}$
@@ 2.3 Limitations and extensions
Third, we have ignored the self-gravity of the perturbed stellar distribution. As shown by Darling & Widrow (2019) and Widrow (2023), self-gravity may strengthen the amplitude of the Snail-like response to an impulsive perturbation, and extend the lifetime of a Snail well beyond that expected due to phase mixing alone. It is unclear what impact self-gravity would have on galactic echoes. On the one hand, we expect that if self-gravity leads to stronger first-order responses, it should lead to stronger second- and higher-order effects like echoes too. On the other hand, the $t_{2}-t_{1}$ dependence of the echo term arises precisely because nonlinear couplings feed off phase-space gradients, and under normal phase mixing these get sharper as $t_{2}-t_{1}$ increases. Since self-gravity will reduce the rate of phase mixing, it may make echoes weaker, not stronger. In a homogeneous plasma, one can perform a self-consistent calculation (Gould et al., 1967; O’Neil & Gould, 1968), which predicts that collective effects amplify the echo while slightly modifying its growth and decay rates. However, to perform such a self-consistent calculation here would incur a significant mathematical penalty: not only would we need to Laplace transform from the time-domain into the complex frequency domain (Gould et al., 1967), but also, since we are dealing here with an inhomogeneous system, we would need to project our fluctuations onto a biorthogonal potential-density basis (e.g., Kalnajs, 1976; Weinberg, 1991; Hamilton & Fouvry, 2024). Alternatively, we could forego both of these transforms and just solve the linearized equations numerically (Widrow, 2023). Since our aim here is to demonstrate the existence of an echo phenomenon rather than perform detailed modeling of the Solar neighborhood, we will not investigate the role of self-gravity further in this paper.
@@ 3.1 Model
We subject our disc to two impulsive perturbations caused by passing satellites or dark-matter subhalos. According to the linear response analysis by Banik et al. (2022, 2023), such encounters can excite both dipole ($n=1$) and quadrupole ($n=2$) modes in the vertical phase-space distribution of stars. The relative amplitudes of these modes depend critically on the parameters of the disc-satellite encounter, in particular, the encounter speed, $v_{\rm p}$, and the angle at which the satellite crosses the disc, $\vartheta_{\rm p}$, where $\vartheta_{\rm p}=0$ corresponds to a perpendicular passage through the disc.
To illustrate how two perturbations with different wave numbers can generate an echo, we adopt a simple scenario in which the first satellite imparts a uniform kick $\Delta v_{1}(z)=\Delta V_{1}$ to stars at $t=t_{1}$, exciting primarily a dipole-like ($n_{1}=1$) perturbation. As a rough estimate, a satellite of mass $M_{\rm p}=4\times 10^{9}\,{\rm M}_{\odot}$, velocity $v_{\rm p}=300\,{\rm kpc}\,{\rm Gyr}^{-1}$, and impact parameter $b=10\,{\rm kpc}$ would impart a kick of order $\Delta V_{1}\sim{\rm G}M_{\rm p}/bv_{\rm p}\sim 6\,{\rm kpc}\,{\rm Gyr}^{-1}$ (see Binney & Schönrich 2018 for a more elaborate modelling). Later, at $t=t_{2}$, we assume that the disc encounters another satellite/subhalo which crosses near the Solar neighborhood at a grazing angle, thereby inducing a quadrupole-like ($n_{2}=2$) perturbation. We determine the $z$ dependence of the second kick by assuming that the perturber is a Plummer sphere, crossing the disc at $\vartheta_{\rm p}=\pi/2$. The net kick imparted to stars in the impulsive limit is then (see Appendix A)
@@
and $a$ is the core radius. The validity of the impulse approximation is examined in Appendix A. In our fiducial setup, we adopt $t_{1}=0.2\,{\rm Gyr}$, $t_{2}=0.8\,{\rm Gyr}$, $\Delta V_{1}=6\,{\rm kpc}\,{\rm Gyr}^{-1}$, $\Delta V_{2}=6\,{\rm kpc}\,{\rm Gyr}^{-1}$ and $a=0.3\,{\rm kpc}$. The corresponding mass of the second satellite/subhalo is $M_{\rm p}\sim 1.2\times 10^{8}\,{\rm M}_{\odot}$, although this estimate can be much larger if the satellite passes at a non-zero distance from the Solar neighborhood.
@@ 3.2 Emergence of echo in the disc’s vertical phase space
The first uniform kick (blue) shifts the entire distribution towards higher velocity, resulting in an overdensity at $\theta=0$ and underdensity at $\theta=\pm\pi$. This dipole ($n_{1}=1$) perturbation subsequently shears into a one-armed phase spiral in the $(z,v)$ plane. In the $(\theta,\Omega)$ plane, this spiral appears as an inclined, rectilinear stripe. The second antisymmetric kick (green) introduces a quadrupole ($n_{2}=2$) perturbation, which evolves into a two-armed phase spiral. In addition to generating its own response, the second kick also perturbs the pre-existing one-armed phase spiral. This second-order perturbation results in a new anti-wound spiral pattern, which appears as a linear pattern with negative inclination in the $(\theta,\Omega)$ plane. As time progresses, this pattern unwinds (phase ‘unmixes’), and at the predicted echo time (red), its phase fully aligns, giving rise to an echo. The echo subsequently phase mixes and forms a new one-armed phase spiral pattern.
The phase alignment at $t_{\rm echo}$ implies that the echo temporarily causes a fluctuation in macroscopic (integrated) quantities. To illustrate this, we plot in the top panel of Fig. 2 the time evolution of the mean height of the disc, $\langle z\rangle=\int{\rm d}z{\rm d}vf(z,v,t)z$. Following the first kick, the mean height undergoes a rapid oscillation, whose amplitude gradually decays due to phase mixing. The second kick induces a density perturbation symmetric about the disc mid-plane and therefore leaves the mean height unaffected. Later, as the echo unwinds and rewinds, the disc’s mean height begins to oscillate again spontaneously, despite the lack of any external forcing. This demonstrates how the memory of earlier perturbations can be preserved in phase space and reappear later in macroscopic space. The remaining panels of Fig. 2 show a similar (though less dramatic) echo effect in the measurement of the disc’s ‘thickness’ $\langle z^{2}\rangle-\langle z\rangle^{2}]^{1/2}$ and in $\langle z^{2}\rangle$ alone. The effect is less pronounced because the echo has most of its power at $n=\pm 1$, which does not contribute to even moments of $z$.
Figure 2: Time evolution of the mean height of the disc $\langle z\rangle$, its thickness $\langle z^{2}\rangle-\langle z\rangle^{2}]^{1/2}$, and the rms vertical coordinate $\langle z^{2}\rangle$, all in the absence of diffusion. The mean height undergoes damped oscillation after the first uniform kick at $t_{1}$ due to phase mixing. The second antisymmetric kick at $t_{2}$ does not directly affect the mean height, although it distorts the pre-existing one-armed phase spiral, generating a renewed oscillation (an echo) near $t_{\rm echo}$. The echo signals in the other panels are weaker (note the shrunken vertical axis).
@@ 4.1 Implications for the Gaia Snail
Although our model is oversimplified in many respects, our simulation shows that two Sagittarius-like kicks can, in principle, generate an echo with amplitudes comparable to those of the Gaia Snails (A=0.1A=0.10.20.2; see Frankel et al. 2023). However, attributing the Snails we observe today entirely to echoes imposes rather stringent requirements.
Suppose, quite reasonably, that the first kick was provided by the Sagittarius dwarf galaxy, which is expected to primarily excite an $n_{1}=1$ mode (e.g., Banik et al., 2023). Fig. 5 then implies that, for the echo to account for the Gaia Snail, the second satellite/subhalo must have passed near the Solar neighborhood at a glancing angle $\sim 0.4\,{\rm Gyr}$ after the passage of Sagittarius and delivered an antisymmetric kick of strength $\Delta V_{2}\gtrsim 6\,{\rm kpc}\,{\rm Gyr}^{-1}$, which requires a mass of $M_{\rm p}\gtrsim 10^{8}\,{\rm M}_{\odot}$3.1, see also Appendix A for the effect of time dependence, which further increases the required mass). However, there are no known luminous satellites of such mass near the Galactic disc apart from Sagittarius (see Banik et al. 2022 and references therein). It is possible that the second kick was caused by an unseen dark subhalo, although such a massive subhalo would likely carry a detectable number of baryons. Moreover, several $N$-body simulations have shown that it is difficult to reproduce the Snail even by Sagittarius alone unless its final mass is much higher than is observed (e.g., Bland-Hawthorn & Tepper-García, 2021; Bennett et al., 2022; Asano et al., 2025). Since the (second-order) echo is inherently weaker than the (first-order) response probed in those studies, we expect that echoes would face even greater difficulty in reproducing the observed amplitude.
A more realistic scenario would be that weak echoes from perturbations by Sagittarius and other satellites/subhalos are continually overlapping with the direct first-order responses, forming phase spirals with complicated variation in amplitude and winding time across both the vertical and horizontal phase space. The residuals between the data and models based on a single encounter indeed show coherent structures, indicating the presence of multiple overlapping phase-spirals (Frankel et al., 2023). Disentangling the precise origin(s) of the Snails remains a challenging task and lies far beyond the scope of this paper. On this front, various stellar labels like age and chemical abundances, which also exhibit Snail-like patterns, may offer valuable clues, as different labels have different gradients in vertical action and angular momentum, and can thus help constrain the nature of the perturbations (Frankel et al., 2025).